A local bank is offering a zero-coupon certificate of deposit for​ $25,000. At​ maturity, three years from​ now, the investor will receive​ $32,000. What is the rate of return on this​ investment?
Answer by Theo(13342)

f = p * (1 + r) ^ n

f = 32 thousand
p = 25 thousand
n = 3 years
r is what you want to find.

equation becomes 32 = 25 * (1 + r) ^ 3

divide both sides of this equation by 25 to get 32 / 25 = (1 + r) ^ 3

take the cube root of both sides of this equation to get (32 / 25) ^ (1/3) = 1 + r

suvbtract 1 from both sides of this equation to get (32 / 25) ^ (1/3) - 1 = r

solve for r to get r = .085767045.

the annual interest rate is 8.5767045%.